# 1]organization=Helion Energy Inc., city=Everett, WA,country=USA

### v1

@@ 5.1 ZEUS D–T validation and calibration
Detector operated at the radial distance, negative high voltage, and digitizer thresholds as specified above. Data was collected for approximately $60\text{\,}\mathrm{s}$ prior to the shot and $700\text{\,}\mathrm{s}$ following the shot. Figure 3a shows the binned rates $R(t)$ with $7.14\text{\,}\mathrm{s}$, $134.4\text{\,}\mathrm{s}$, and $\geq$400\text{\,}\mathrm{s}$\$ components fitted. A set of reference shots of different yields establish linearity between the Silver detector derived facility fluences and both components of the SiO2 activation detector. The fit-derived $N_{0}^{(^{16}\mathrm{N})}$ and $N_{0}^{(^{28}\mathrm{Al})}$ were plotted and fitted calibration constants of $K_{\mathrm{Al}}=(5.63\pm 0.08)\times 10^{4}$ and $K_{\mathrm{N}}=(4.78\pm 0.07)\times 10^{3}$ (Figure 4). The precise temporal decomposition into $7.13\text{\,}\mathrm{s}$ and $134\text{\,}\mathrm{s}$ components, matching known nuclear data, confirms the signal arises from the claimed activation channels and rules out alternative explanations for the linear trend, such as PMT afterpulsing or slow phosphorescence. In Figure 4, the error bars are derived from the uncertainty of the fitting process.
@@ 5.2 MJOLNIR D–D insensitivity
The detector operated near the MJOLNIR D–D DPF under the same DAQ settings as specified above. By placing the detector closer to the pinch, it was exposed to a nominal D–D fluence comparable to the lower-end ZEUS D–T shots. No count rate increase was observed besides the prompt detection of the shot (Fig. 5). This null result is attributed to neutron energy rather than fluence: the ENDF/B-VII.1 cross sections for $^{28}\mathrm{Si(n,p)}$ and $^{16}\mathrm{O(n,p)}$ are effectively zero at $2.45\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ (Fig. 2), and applying the ZEUS-derived calibration constants shows that $14.1\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ neutrons at this nominal fluence would produce a detectable signal. Dense plasma focus devices operating with deuterium can produce a secondary 14.1 MeV component via burn-up of $\sim$1 MeV tritons created in the D(d,p)T branch. Triton burn-up is a standard feature of deuterium fusion plasmas, with the DT/DD ratio depending on triton confinement and plasma density [23, 24, 25]. In magnetically confined deuterium plasmas, measured burn-up fractions $Y_{DT}/Y_{DD}$ are typically $10^{-3}$ to $10^{-2}$ [25, 26]. For short-pulse DPF plasmas, significantly smaller fractions are expected; values of order $10^{-3}$ to $10^{-4}$ provide a conservative bound. At MJOLNIR fluence levels ($\sim 10^{8}$ to $10^{9}$ n/cm2 at 2.45 MeV), this implies a secondary 14.1 MeV fluence of $\sim 10^{4}$ to $10^{6}$ n/cm2, well below the present detection threshold ($\sim 10^{7}$ n/cm2). The null result is therefore consistent with both the detection threshold and the expected secondary D–T yield.
@@ 6.1 Operating envelope
The detector provides rapid automated yield readout with minimal operator burden. The practical lower yield is set by counting statistics (order $10^{7}$ n/cm2/shot for the present geometry and lowest threshold). At high yields, early-time pileup and list-mode throughput limits can cause digitizer dropouts (Fig. 7); mitigation includes increased standoff, a smaller SiO2 radiator, higher thresholds, or higher-throughput firmware. From Figure 7, the digitizer max rate is observed to be $\sim 10^{4}$ counts/s, as observed by the dropouts in the digitizers sustained output. Raising the digitizer threshold as observed in Figure 6 lowers observed count rate.

### v2

# arXiv:2602.22477v2[physics.ins-det] 27 Feb 2026

@@ 5.1 ZEUS D–T validation and calibration
Detector operated at the radial distance, negative high voltage, and digitizer thresholds as specified above. Data was collected for approximately $60\text{\,}\mathrm{s}$ prior to the shot and $700\text{\,}\mathrm{s}$ following the shot. Figure 3a shows the binned rates $R(t)$ with $7.14\text{\,}\mathrm{s}$, $134.4\text{\,}\mathrm{s}$,ă and $\geq$400\text{\,}\mathrm{s}$\$ components fitted. A set of reference shots of different yields establish linearity between the Silver detector derived facility fluences and both components of the SiO2 activation detector. The fit-derived $N_{0}^{(^{16}\mathrm{N})}$ and $N_{0}^{(^{28}\mathrm{Al})}$ were plotted and fitted calibration constants of $K_{\mathrm{Al}}=(5.63\pm 0.08)\times 10^{-4}\,\text{counts}\cdot\text{cm}^{2}/\text{n}$ and $K_{\mathrm{N}}=(4.78\pm 0.07)\times 10^{-3}\,\text{counts}\cdot\text{cm}^{2}/\text{n}$ (Figure 4). The precise temporal decomposition into $7.13\text{\,}\mathrm{s}$ and $134\text{\,}\mathrm{s}$ components, matching known nuclear data, confirms the signal arises from the claimed activation channels and rules out alternative explanations for the linear trend, such as PMT afterpulsing or slow phosphorescence. In Figure 4, the error bars are derived from the uncertainty of the fitting process.
@@ 5.2 MJOLNIR D–D insensitivity
The detector operated near the MJOLNIR D–D DPF under the same DAQ settings as specified above. By placing the detector closer to the pinch, it was exposed to a nominal D–D fluence comparable to the lower-end ZEUS D–T shots. No count rate increase was observed besides the prompt detection of the shot (Fig. 5). This null result is attributed to neutron energy rather than fluence: the ENDF/B-VII.1 cross sections for $^{28}\mathrm{Si(n,p)}$ and $^{16}\mathrm{O(n,p)}$ are effectively zero at $2.45\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ (Fig. 2), and applying the ZEUS-derived calibration constants shows that $14.1\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ neutrons at this nominal fluence would produce a detectable signal. Dense plasma focus devices operating with deuterium can produce a secondary 14.1 MeV component via burn-up of $\sim$1 MeV tritons created in the D(d,p)T branch. Triton burn-up is a standard feature of deuterium fusion plasmas, with the DT/DD ratio depending on triton confinement and plasma density [23, 24, 25]. In magnetically confined deuterium plasmas, measured burn-up fractions $Y_{DT}/Y_{DD}$ are typically $10^{-3}$ to $10^{-2}$ [25, 26]. For short-pulse DPF plasmas, significantly smaller fractions are expected; values of order $10^{-3}$ to $10^{-4}$ provide a conservative bound. At MJOLNIR fluence levels ($10^{8}$ to $10^{9}$ n/cm2 at 2.45 MeV), this implies a secondary 14.1 MeV fluence of $\leq 10^{4}$ n/cm2. The null result is therefore consistent with both the detection threshold and the expected secondary D–T yield.
@@ 6.1 Operating envelope
The detector provides rapid automated yield readout with minimal operator burden. The practical lower yield is set by counting statistics required for the fit, where fits to the $^{16}\mathrm{N}$ decay have been observed to converge for fluences of $\sim 10^{2}$ n/cm2. At high yields, early-time pileup and list-mode throughput limits can cause digitizer dropouts (Fig. 7); mitigation includes increased standoff, a smaller SiO2 radiator, higher thresholds, or higher-throughput firmware. From Figure 7, the digitizer max rate is observed to be $\sim 10^{4}$ counts/s, as observed by the dropouts in the digitizers sustained output. Raising the digitizer threshold as observed in Figure 6 lowers observed count rate.